OpenStudy (mind2000):

In an illustration of a ladybug, the length of the ladybug is 5.6 centimeters. The actual length of the ladybug is 0.7 centimeter. What is the scale of the drawing?

OpenStudy (lncognlto):

Magnification = measured/actual = 5.6/0.7 =8

OpenStudy (mind2000):

a. 1 cm = 1.25 cm c. 1 cm = 0.125 cm b. 1 cm = 0.125 m d. 1 cm = 12.5 cm

OpenStudy (lncognlto):

Ah, well then the answer is c, 1 cm = 1.25 cm. 5.6/5.6 = 1, 0.7/5.6 = 0.125.

OpenStudy (lncognlto):

Sorry I meant the answer is 1 cm = 0.125 cm.